OpenStudy (anonymous):

Let f(x) = 9x – 2 and g(x) = –x + 3. Find f(g(x)).

OpenStudy (anonymous):

–9x + 27 –9x + 25 –9x + 5 –9x – 2

OpenStudy (anonymous):

f(x) = 9x – 2 f(g(x)) = 9g(x) - 2 but since g(x) = -x + 3, f(g(x)) = 9(-x + 3) - 2 = -9x + 27 - 2 = -9x + 25

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

@brookie5150 YOU'RE THE BEST

OpenStudy (anonymous):

@Reemxoxo , THANKS. hahah